A sample of sodium has a mass Of 46 g. What is the mass Of the same number of calcium atoms as sodium atoms present in given sample ?
1
Expert's answer
2020-11-14T13:51:21-0500

Moles of sodium = 46/23 = 2moles

1 mole of sodium contains 6.023×1023 Atoms

2 moles contains 12.046×10223 atoms


So same number of calcium atoms means 2 moles of calcium.


Mass of calcium = 2×40 = 80 grams